# Geometric Knowledge-Guided Localized Global Distribution Alignment for Federated Learning

### v1

@@ 3.1 Computation of Global Geometric Shapes Under Privacy Constraints
By performing eigenvalue decomposition on $\Sigma_{i}$, we can derive the geometric shape of the global data distribution for class $i$ without sharing any data.
@@ 3.2.2 Multi-Domain Federated Learning
Algorithm 2 GGEUR (Multi-Domain Scenario)
2:$X_{\text{new}}^{i}$: Augmented sample set of class $i$ at client $k$
4:for $j=1$ to $n_{k}^{i}$ do
6:end $\triangleright$ Step 2: Cross-Domain Simulation
7:for each prototype $\mu_{k^{\prime}}^{i}$ from other domains do
9:end for
10:return $X_{\text{new}}^{i}$
@@ 4.2 Evaluation Results on Label Skew Dataset
Main Results. Tables 1, 2, and 3 show the performance improvement of GGEUR over FedAvg (CLIP+MLP) [39] under different $\beta$ values, along with comparison results with other methods. Under all label distribution skew settings, GGEUR significantly outperforms other methods, demonstrating superior classification accuracy and adaptability to imbalanced label distributions. Particularly at lower $\beta$ values (i.e., with more severe label skew), GGEUR notably enhances the performance of FedAvg (CLIP+MLP) on CIFAR-100, validating its effectiveness in handling extreme label skew scenarios. For example, when $\beta$ is $0.01$, $0.03$, and $0.05$, GGEUR achieves performance gains of 17.01%, 14.63%, and 13.64%, respectively. These results indicate that GGEUR not only exhibits good generalization ability in standard settings but also maintains robust performance in cases of significant label skew.
@@ 4.3 Evaluation Results on Domain Skew Dataset
Main Results. Tables 4, 5, and 6 present the experimental results on datasets with domain skews. It can be observed that simply using CLIP [47] for image representation, combined with a single-layer MLP for federated learning, already surpasses existing methods. This improvement is attributed to the advancements in the foundation model, and GGEUR can further significantly enhance the performance of the global model. For instance, on the Digits dataset, GGEUR improves the average performance of FedAvg (CLIP+MLP) [39], MOON (CLIP+MLP) [25], and FedProto (CLIP+MLP) [55] by 5.76%, 18.06%, and 5.38%, respectively. Additionally, compared to other methods, GGEUR significantly reduces the accuracy variance across different domains. This demonstrates that GGEUR effectively adapts to features from different domains when handling cross-domain data distribution disparities, providing more robust and fair performance.
@@ Appendix B Dataset
In the Office-Home dataset, while ensuring that each client corresponds to a single domain, we first generate a Dirichlet coefficient matrix, where the degree of class imbalance is controlled by $\beta$. For the $65$-class, $4$-domain Office-Home task, we generate a $4\times 65$ matrix controlled by $\beta$ (with each column summing to $1$). The four coefficients for each class are then allocated to the four clients, and each client uses its assigned coefficients to determine the number of samples for that class. This setup results in a distribution that incorporates both domain shift (one domain per client) and Dirichlet-based class imbalance, presenting a scenario where the model faces both class distribution and domain differences, creating a more realistic, challenging, and diverse distribution for classification. We name the newly constructed dataset Office-Home-LDS (Label and Domain Skew). Figure 9 shows the data distribution of Office-Home-LDS with different $\beta$ values.

### v2

@@ 3.1 Computation of Global Geometric Shapes Under Privacy Constraints
where each client’s contribution to $\Sigma_{i}$ is determined by the coefficient $\frac{N_{k}^{i}}{N_{i}}$, By performing eigenvalue decomposition on $\Sigma_{i}$, we can derive the geometric shape of the global data distribution for class $i$ without sharing any data. For more details on privacy constraints, please refer to Appendix D.
@@ 4.2 Evaluation Results on Label Skew Dataset
Main Results. Tables 1, 2, and 3 show the performance improvement of GGEUR over FedAvg (CLIP+MLP) [39] under different $\beta$ values, along with comparison results with other methods. Under all label distribution skew settings, GGEUR significantly outperforms other methods, demonstrating superior classification accuracy and adaptability to imbalanced label distributions. Particularly at lower $\beta$ values (i.e., with more severe label skew), GGEUR notably enhances the performance of FedAvg (CLIP+MLP) on CIFAR-100, validating its effectiveness in handling extreme label skew scenarios. For example, when $\beta$ is $0.01$, $0.03$, and $0.05$, GGEUR achieves performance gains of 17.01%, 14.63%, and 13.64%, respectively. These results indicate that GGEUR not only exhibits good generalization ability in standard settings but also maintains robust performance in cases of significant label skew. Moreover, our large-scale client experiments further confirm the scalability and consistent effectiveness of GGEUR under increased client participation, details are provided in Appendix E.
@@ 4.3 Evaluation Results on Domain Skew Dataset
Main Results. Tables 4, 5, and 6 present the experimental results on datasets with domain skews. It can be observed that simply using CLIP [47] for image representation, combined with a single-layer MLP for federated learning, already surpasses existing methods. This improvement is attributed to the advancements in the foundation model, and GGEUR can further significantly enhance the performance of the global model. For instance, on the Digits dataset, GGEUR improves the average performance of FedAvg (CLIP+MLP) [39], MOON (CLIP+MLP) [25], and FedProto (CLIP+MLP) [55] by 5.76%, 18.06%, and 5.38%, respectively. Additionally, compared to other methods, GGEUR significantly reduces the accuracy variance across different domains. This demonstrates that GGEUR effectively adapts to features from different domains when handling cross-domain data distribution disparities, providing more robust and fair performance. We also conducted experiments to measure the computational cost of GGEUR, which demonstrated that the additional overhead introduced by our method is minimal. Further details can be found in Appendix F.
@@ Appendix B Dataset
In the Office-Home dataset, while ensuring that each client corresponds to a single domain, we first generate a Dirichlet coefficient matrix, where the degree of class imbalance is controlled by $\beta$. For the $65$-class, $4$-domain Office-Home task, we generate a $4\times 65$ matrix controlled by $\beta$ (with each column summing to $1$). The four coefficients for each class are then allocated to the four clients, and each client uses its assigned coefficients to determine the number of samples for that class. This setup results in a distribution that incorporates both domain shift (one domain per client) and Dirichlet-based class imbalance, presenting a scenario where the model faces both class distribution and domain differences, creating a more realistic, challenging, and diverse distribution for classification. We name the newly constructed dataset Office-Home-LDS (Label and Domain Skew). Figure 9 shows the data distribution of Office-Home-LDS with different $\beta$ values. Dataset and Constructor published at: https://huggingface.co/datasets/WeiDai-David/Office-Home-LDS.
@@ Appendix C Implementation Details
Appendix D Privacy Constraints
In our approach, the server only sends the eigenvectors and eigenvalues of the global covariance matrix back to clients, without sharing raw data or local covariance matrices. We demonstrate below that this information is insufficient for reconstructing any client’s original data.
(1) Eigenvectors and Eigenvalues Do Not Contain Raw Data. The eigendecomposition provides only the geometric structure of the data distribution, without encoding individual sample details. Even if a client obtains eigenvectors and eigenvalues, reconstructing the original data is an ill-posed problem, as it admits infinitely many solutions.
(2) Low-Rank Property Prevents Data Reconstruction. The covariance matrix is typically low-rank, meaning: $\text{rank}(\Sigma_{i})\ll d$, where $d$ is the original data dimension. This implies that even with full knowledge of eigenvectors and eigenvalues, clients can only access principal directions of the data and not its full details.
(3) Aggregation Prevents Isolation of Individual Client Contributions. The global covariance matrix is an aggregate of all clients’ local covariance matrices: $\Sigma_{i}=\sum_{k=1}^{K}\frac{n_{k}^{i}}{N_{i}}\Sigma_{k}^{i}+\sum_{k=1}^{K}\frac{n_{k}^{i}}{N_{i}}(\mu_{k}^{i}-\mu_{i})(\mu_{k}^{i}-\mu_{i})^{T}.$ Since each client’s contribution is mixed through weighted averaging: I. No single client can isolate another client’s contribution from Geometric Knowledge. II. Even if a client’s data is removed, the impact on $\Sigma_{i}$ is distributed across all eigenvectors, making individual influences indistinguishable.
(4) Existing Literature Supports the Privacy of Covariance Matrices. Prior works confirm that sharing higher-order statistics (covariance matrices) poses lower privacy risks than sharing model gradients (Melis et al.).
Appendix E Large-Scale Client
In practical federated learning settings, the number of participating clients can significantly affect model performance due to increased data heterogeneity. To further evaluate the performance of our proposed method under larger-scale federated learning scenarios, we conducted additional experiments with an increased number of clients. Specifically, we conducted experiments on the label-skewed dataset CIFAR-10 with 100, 300, and 500 clients. As shown in Table 10, the results demonstrate that GGEUR remains robust and continues to enhance the performance of FedAvg (CLIP+MLP).
Table 10: Number of Clients $K$ Impact on Performance.
  CIFAR-10 ($\beta=0.1$)
Methods $K=100$ $K=300$ $K=500$
FedAvg (CLIP+MLP) 87.89 84.69 82.05
+ GGEUR 93.55 ($+5.66$) 92.17 ($+7.84$) 90.43 ($+8.38$)
Appendix F Computational Cost
We conducted experiments on the domain-skewed dataset Digits, comparing the training time required to complete the full model for FedAvg (CLIP+MLP), SCAFFOLD (CLIP+MLP), and MOON (CLIP+MLP) before and after applying GGEUR. The results (Tab11) show that GGEUR introduces almost no additional training time overhead. Specifically, after applying GGEUR, the training time for the three methods increased by only 3.5s, 4.6s, and 3.3s, respectively.
Table 11: The average training time (s) per round.
  Digits
Methods FedAvg SCAFFOLD MOON
CLIP+MLP 28.2 54.5 32.3
+ GGEUR 31.7 (+3.5) 59.1 (+4.6) 35.6 (+3.3)
Algorithm 2 GGEUR (Multi-Domain Scenario)
2:$X_{\text{new}}^{i}$: Augmented sample set of class $i$ at client $k$
4:for $j=1$ to $n_{k}^{i}$ do
6:end $\triangleright$ Step 2: Cross-Domain Simulation
7:for each prototype $\mu_{k^{\prime}}^{i}$ from other domains do
9:end for
10:return $X_{\text{new}}^{i}$